Calculate the mean deviation and it's coefficient from median for the following.
X: 12 17 18 19 17 16 17 13
1
Expert's answer
2019-10-04T11:01:22-0400

Arranged array: 12, 13, 16, 17, 17, 17, 18, 19.

"median=\frac{17+17}{2}=17."

"M.D.=\frac{1}{n}\sum|x-median|=\frac{1}{8}(5+4+1+0+0+0+1+2)=\\=\frac{13}{8}=1.625."

"Coefficient\; of \;M.D.=\frac{M.D.}{median}=\frac{13}{136}=0.0956."